OpenStudy (igster1112):

Simplify: 3 1/3 - 2/5

OpenStudy (anonymous):

2.93 aprox

OpenStudy (sepeario):

Change them into improper fractions first; 10/3 - 2/5 =

OpenStudy (agl202):

2 and 14/15 is actual answer!

OpenStudy (xapproachesinfinity):

3 +1/3-2/5 =?

OpenStudy (igster1112):

Ok I changed it into an improper graction.

OpenStudy (sepeario):

Then find the common denominator and you get 50/15 - 6/15 = 44/15 or 2 and 14/15

OpenStudy (xapproachesinfinity):

3 1/3 is the same as 3+1/3

OpenStudy (xapproachesinfinity):

you don't have to change anything if they didn't tell you to give answer in improper fraction

OpenStudy (sepeario):

Haha good point @xapproachesinfinity but the answer isn't 3 and 1/3

OpenStudy (igster1112):

It says simplify.

OpenStudy (xapproachesinfinity):

i didn't say the answer is 3 and 1/3 i wrote 3+1/3-2/5 @Sepeario

OpenStudy (sepeario):

Oh sorry

OpenStudy (xapproachesinfinity):

no worries!